OpenStudy (anonymous):

Three times the square of a nonzero number is equal to twelve times the number What is the number?

OpenStudy (anonymous):

3*x^2=12*x

OpenStudy (anonymous):

do you understand the equation @Kristen_D17

OpenStudy (anonymous):

you basically just plug in random numbers starting at 1 until you get your answer

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

1 doesn't work because the square of 1 is just one. and 1 time 12 i twelve which =/= 1

OpenStudy (anonymous):

3x^2=12x. divide the x on bothsides to get 3x=12

OpenStudy (anonymous):

2 doesn't work and neither does 3. try four

OpenStudy (anonymous):

you should end up with 4 as your answer

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

try to break the question down as much as possible before trying to solve the equation.

OpenStudy (anonymous):

you could also just solved the equation i provided for you.

OpenStudy (anonymous):

but sometimes, guessing and plugging and chugging is a good method to go with, but sometimes its just faster to solve the equation.

OpenStudy (anonymous):

which ever is easier to use is up to you

OpenStudy (anonymous):

right because sometimes when your trying to plug and chug, the answer could be in the thousands and that would take forever.